OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form?

OpenStudy (anonymous):

y=x-6

OpenStudy (anonymous):

x + y = –6 x – y = –6 x – y = 6 x + y = 6 are the possible answers

OpenStudy (anonymous):

x+y=-6

OpenStudy (anonymous):

thanks for the medal :)

OpenStudy (anonymous):

It's x-y=6

OpenStudy (anonymous):

y=ax+b parallel --> a=1 y=x+b through (-3,-9) -9=-3+b --> b=-6 y=x-6 x-y=6

hero (hero):

:/ I had already posted the solution